OpenStudy (anonymous):

can you help me to find the LCM? I am not sure how to do this one.. Thank you w^3+10w^2+25w, w^2-10w

OpenStudy (a_clan):

factorize both

OpenStudy (anonymous):

LCM is mean find bigest number w, w^2=w*w w^3=w*w*w 10=2*5 25=5*5 LCM=5*5*2 =5 than LCM= 50w^3

OpenStudy (oblivious):

First from all we factor out w (w^2 +10 w +25)*w and (w - 10)*w then (w+5)^2 *w and w -10*w so w = -5 in the first case and -10 in the second (unless w = 0, which is an irrelevent case) the LCM of -5 and -10 is, of course, -10

OpenStudy (anonymous):

Thank you, I am still having a little trouble, but I am sure with practice I will understand it better